OpenStudy (anonymous):

Which test score is greater than 80 %? 12 out of 20 35 out of 45 11 out of 15 33 out of 40

OpenStudy (anonymous):

well 12/20 = 3/5 = 0.6 = 60% and 35/45 = 7/9 = 0.777777... = 77.77% and 11/15 = 0.7333333333333333333333..... = 73.33% and 33/40 = 0.825 = 82.5% so D

OpenStudy (anonymous):

Thanks! @Allofus